Hi there,
How do you calculate the work done by friction given the coefficient of friction, mass, incline angle and distance?

Problem: A 12.2 kg mass Is moving down a rough incline plane with an initial speed Vo. The incline has a coefficient friction of 0.596 and is inclined at an angle of 33 degrees horizontally. The mass moves a distance of 2.77 m along a plane before stopping. calculate the work done by friction (in J).

Any info would be appreciated.
1
Expert's answer
2020-11-16T07:49:34-0500


Here "\mathbf{F}_{fr}" is frictional force, "\mathbf{N}" is the normal force and "m\mathbf{g}" is gravitational force. "\alpha = 33\degree" is the angle of inclination.

By definition, "F_{fr} = \mu N", where "\mu = 0.596".

Let's write down Newton's second law in projection on y-axis.


"N - mg\cos\alpha = 0\\
N =mg\cos\alpha"

Subsitiuting the last expression into the definition of "F_{fr}" , obtain:


"F_{fr} = \mu N = \mu mg\cos\alpha"

The work done by friction is given then as


"A = F_{fr}\cdot d = d\mu mg\cos\alpha"

where "d = 2.77m". Thus, obtain:


"A = 2.77\cdot 0.596 \cdot 12.2\cdot 9.81\cdot \cos33\degree \approx 165.7\space J"

Answer. 165.7 J.